### v1

# Dynamic Q&A of Clinical Documents with Large Language Models

Abstract
Electronic health records (EHRs) contain a wealth of valuable patient information embedded within clinical notes. However, the ever-increasing volume and complexity of these unstructured clinical notes makes it extremely challenging for researchers and clinicians to manually locate and extract relevant information in a timely manner. Intelligent systems that can understand clinical documents and enable users to efficiently query and retrieve key details on demand are urgently needed. In this work our aim is to develop a natural language conversational interface powered by state-of-the-art open source large language models (LLMs) that allows users to explore clinical notes through dynamic question-answering. A chatbot interface was built using Langchain, a framework for creating and managing applications powered by large language models, paired with powerful transformer-based language models. The system allows users to simply ask questions in natural language and receive answers extracted from relevant parts of the clinical notes. Extensive experiments were performed to evaluate combinations of semantic embedding models such as the state of the art SentenceTransformers framework, based on the Sentence-BERT embedding models [1], and advanced large language models on their ability to encode both queries and documents for optimal information retrieval. Wizard Vicuna demonstrated the highest accuracy due to its 13B parameters but required heavy compute. To improve inference latency and deployability, model optimization techniques like weight quantization was utilized to reduce latency by approximately 48 times. While results are promising, several limitations remain including model hallucinations, and lack of robust evaluation across diverse medical cases. Addressing these gaps represents important future work to unlock the value in clinical notes and advance clinical decision making through AI.
1 Introduction
Electronic health record (EHR) systems have become ubiquitous in modern healthcare environments [2, 3, 4]. They offer an integrated digital repository of patient health information collected from various clinical encounters across time. EHRs may contain a range of data including demographics, medical and family histories, laboratory test results, radiology images, computed biosignals and crucially, clinical notes generated by physicians, nurses and other providers. These clinical notes, documenting physician-patient interactions, encapsulate a wealth of vital information about the patient’s conditions, treatments, care plans and outcomes in rich natural language. However, a significant limitation of EHRs is that critical patient information is buried within large volumes of unstructured clinical notes that are challenging for providers to efficiently search through and extract relevant details on demand. For instance, a newly assigned physician may need to rapidly comprehend a patient’s care history across multiple specialists to make appropriate decisions. Or a researcher may want to analyze treatment response patterns across populations. These use cases require intelligently sifting through troves of notes to pinpoint key facts, relationships and temporal trends - a formidable manual task. Hence, there is a profound need for intelligent systems that can automatically understand clinical notes written in natural language and enable users to easily query and retrieve targeted information on the fly through conversational interactions.
Traditionally, information retrieval systems have relied on methods like Boolean models, probabilistic models, and term frequency-inverse document frequency (TF-IDF) for surface-level matching of keywords between queries and documents. While these techniques have been useful for basic keyword-based retrieval, they have their limitations. Boolean models utilize operators like AND, OR, and NOT to retrieve documents containing specific keyword combinations, but they lack the ability to understand the nuances and context of natural language queries. Probabilistic models rank results based on keyword occurrences, which can be effective for certain tasks, but they struggle to grasp the underlying meaning of the text. TF-IDF is another widely used method that assigns weights to keywords in queries and documents based on their rarity. This approach helps identify important terms, but it still falls short in capturing the full semantic context and relationships between words. However the primary constraint of these methods is that they lack semantic understanding and the ability to interpret context.
As a result, these traditional techniques often miss relevant information when applied to complex natural language text. They fail to recognize synonyms, antonyms, word variations, and other language intricacies that are essential for accurate and comprehensive information retrieval.
To overcome these limitations and unlock deeper meaning, more advanced techniques such as natural language methods based on large language models (LLMs) [5] can be utilized. The rapidly advancing LLMs, like GPT-3 and its successors, have revolutionized the field of natural language processing, with their contextual understanding, generative capabilities, and fine-tuning adaptability. Integrating LLMs into information retrieval systems holds the potential to deliver more accurate, insightful, and context-aware results, transforming the way users access information in a more personalized and efficient manner.
In this project we explore the utility of modern natural language processing (NLP) methods based on large language models (LLMs) to develop a conversational agent or chatbot that allows users to explore clinical notes through dynamic question answering. By interfacing with the chatbot in natural language as one would with a human assistant, users can extract relevant details buried within notes, facilitating improved clinical care and research.
Contributions: To this end, our key contributions are as follows:
  • We developed a conversational question answering system based large language models that allows natural language interaction for exploring clinical notes
  • We discussed detailed step by step pipeline of the question answering system based on a medical dataset
  • We conducted extensive experiments evaluating combinations of different embedding models and LLMs to determine optimal configurations for accuracy and speed
  • We showed how the latency and deployability of these models can be optimized by using weights quantization
  • We highlighted the superiority of retrieval augmented generation over domain-specific fine-tuning, demonstrating its advantages in our study
  • We highlighted key limitations around model performance, evaluation rigor, and real-world deployment that open the doors for future research in that domain
Organization: The paper is organized as follows: We discuss related works in section 2. Section 3 discusses the approach. Details of the dataset are discussed in section 4. Evaluation and results are discussed in section 5. We discuss the limitations and future work in section 6 and section 7 concludes the paper.
2 Related Work
2.1 LLMs in information retrieval
The intersection of Large Language Models (LLMs) and Information Retrieval (IR) has emerged as a pivotal area of research, transforming traditional approaches and reshaping the landscape of information access.
The work by Zhu et al. [6] provides a strong foundation by offering a comprehensive survey of the diverse applications of LLMs in different facets of IR, including query rewriting, retrieval, re-ranking, and reading. While their survey spans a wide array of studies, our research takes a specific angle, honing in on the integration of LLMs in the information retrieval process. Expanding upon this foundation, we delve into the potential of enhancing traditional retrieval methods by incorporating text generation capabilities. Our aim is to provide a more natural and conversational interaction between users and the system, ensuring a seamless and user-friendly experience.
Another influential contribution in the realm of information retrieval is the tutorial presented by Begoli et al. [7]. This tutorial focuses on advancing both the theoretical and practical aspects of transformer-based approaches, particularly with an emphasis on efficiency and robustness in large-scale information retrieval scenarios. The tutorial aims to equip the audience with insights gained from developing and fine-tuning transformer/BERT-based models for IR tasks. In contrast, our approach focuses on the utilization of off-the-shelf models, enabling the direct retrieval of information from provided resources without the need for intricate fine-tuning procedures. This streamlined method significantly boosts usability and efficiency, circumventing the time and computational expenses associated with fine-tuning. Furthermore, our research demonstrates that the accuracy achieved through our direct utilization method surpasses that attained through domain-specific fine-tuning processes.
2.2 Medical question answering using LLMs
In recent years, the field of medical question answering has witnessed remarkable advancements, driven by the integration of Large Language Models (LLMs) into the domain. The quest for expert-level performance in this grand challenge has seen notable contributions from various research works.
Lee et al. [8] address the unique challenges posed by biomedical text mining. Recognizing the limitations of applying general NLP advancements to biomedical corpora, they introduce BioBERT, a domain-specific language representation model pre-trained on large-scale biomedical corpora. BioBERT outperforms generic models like BERT and previous state-of-the-art models in various biomedical text mining tasks, including biomedical named entity recognition, biomedical relation extraction, and biomedical question answering.
Another work by Chakraborty et al. [9]emphasizes the urgency of accelerating biomedical research in light of the SARS-CoV-2 pandemic. Introducing BioMedBERT, a pre-trained model on the extensive BREATHE v1.0 dataset, the authors showcase its efficacy in Question-Answering (QA) and Information Retrieval (IR) tasks. The model achieves state-of-the-art results on BioASQ datasets and exhibits superior performance when coupled with Elasticsearch for IR tasks on the BioASQ dataset.
Singhal et al. [10] present Med-PaLM 2, a groundbreaking development in medical question answering. Addressing the inherent challenges in prior models, Med-PaLM 2 leverages a combination of improvements in base LLMs (specifically PaLM 2), medical domain fine-tuning, and innovative prompting strategies, including a novel ensemble refinement approach. The model achieved an impressive score of up to 86.5% on the MedQA dataset, surpassing its predecessor by over 19% and establishing a new state-of-the-art.
In contrast to the aforementioned works, our work diverges by placing a specific emphasis on the integration of retrieval augmented generation for medical question answering. While the discussed models, such as Med-PaLM 2, BioBERT, and BioMedBERT, have made significant strides in achieving expert-level performance through advancements in base LLMs, domain fine-tuning, and innovative prompting strategies, my approach focuses on enhancing deployability and mitigating computational and time constraints. The integration of retrieval augmented generation aims to strike a balance between performance and resource efficiency, making the model more accessible for widespread adoption in medical settings.
2.3 Impact of embedding models on output of large language models
The exploration of word embeddings and their influence on the output of large language models (LLMs) is an under-explored but crucial aspect in natural language processing (NLP). While numerous studies focus on the applications and performance of LLMs, the specific impact of embedding models on their output remains relatively uncharted territory.
The only relevant work in this area is the work by Asudani et al. [11]. In this work the authors focuses on the application of word embeddings and deep learning models in the context of text analytics tasks. Through a systematic literature review, the authors classify and analyze relevant articles, highlighting the increasing interest in employing these techniques for analysis and prediction. The study categorizes literature based on text analytics applications, encompassing areas like text classification, sentiment analysis, named entity recognition (NER), recommendation systems, biomedical text mining, and topic modeling. The paper concludes that utilizing domain-specific word embeddings and the LSTM model can significantly enhance overall text analytics task performance. In our investigation, we direct our attention to a distinct domain that remains unexplored in the aforementioned study — medical question answering. Furthermore, we delve into a different suite of embedding models and language models, areas that Asudani et al. did not specifically address in their research.
3 Approach
Figure 1: Information retrieval pipeline
Our approach primarily utilizes Retrieval-Augmented Generation (RAG) [12], a technique that merges the power of pre-trained language models with the benefits of external knowledge sources. RAG employs a two-step process: a retrieval step that fetches relevant documents from a vast corpus of text, and a generation step that crafts responses based on these retrieved documents and the given input query.
This technique is particularly advantageous when the required response includes information that the language model has not encountered during its training. By combining the ability to generalize from language model pretraining with the capacity to retrieve specific details from external sources, RAG can generate highly detailed and accurate responses while avoiding the overhead of fine tuning the language model on specific datasets.
As illustrated in Figure 1 our approach consists of two main phases: a) Ingestion b) Generation.
  1. 1.
    Ingestion
    • Content Extraction: We begin by extracting the content of the clinical documents from the objects that encapsulate both the content and metadata (i.e source path, file size,…etc).
    • Content Splitting: The loaded content is then split into chucks. This step is crucial as it breaks down the text into smaller units, to comply with the embedding models context size limitations. Throughout this step, we intentionally allow for some overlap between the different chunks. This strategy helps to maintain the contextual integrity of the data.
    • Text Embedding: Each chunk of text is then transformed into an embedding. These embeddings are numerical representations that encapsulate the semantic meaning of the text. The transformation process is carried out by a distinct embedding model, the details of which will be discussed later.
    • Vector Store: The embeddings, along with some reference to the original content the embedding was created from, are subsequently stored in a vector database or vector store, which forms the backbone of our retrieval system.
  2. 2.
    Generation
    • Document Retrieval: The user input query is processed using the same process described above 1. After that, an index – a data structure designed for quick retrieval – is used by a retriever component to scan the vector store. Using a cosine similarity, the retriever identifies and returns the most relevant documents to the user’s input query.
    • Output Generation: The re-ranked vectors, which represent the selected documents, along with the input prompt, are passed to a language model. This model is responsible for text summarization and final output generation, and it is distinct from the earlier mentioned embedding model.
The embedding model and the language model are two separate but integral components of our system. The embedding model is used to transform raw text into meaningful numerical representations, while the language model generates coherent and contextually appropriate text. A summary of the distinguish between embedding and language models is shown in table 1
Model type Description Applications
Language Models (LMs)
Understand and generate human-like text,
learning patterns from large data
•Text completion
•Language generation
Embedding Models
Encode text into vector representations called embeddings,
capturing semantic meaning.
•Comparison and similarity calculations
Table 1: Embedding vs Language models
To achieve the previously described pipeline we utilized a tool called Langchain. LangChain is a powerful and versatile Python framework designed to facilitate the development of state-of-the-art language models and their applications. It offers a wide range of pre-configured models and utilities that allow researchers and developers to focus on the creative aspects of their projects, rather than spending time on repetitive and time-consuming tasks such as model configuration, training, and deployment. With its modular and flexible design, LangChain allows users to easily integrate various components and adapt them to their specific needs.
In our approach, we leveraged multiple components of the LangChain framework. First, we used the Document Store component for efficiently storing and retrieving documents. This component allowed us to create a robust retrieval system, which is a crucial aspect of our RAG model. The Document Store uses a vector database to store embeddings and references to the original content, providing a fast and scalable solution for storing and retrieving large volumes of data.
Second, we employed LangChain’s Retriever component for document retrieval. The retriever uses an index, a data structure designed for quick retrieval, to scan the vector store. Using cosine similarity, it identifies and returns the most relevant documents to the user’s input query. This component’s ability to quickly and accurately retrieve relevant documents is essential for the successful application of our RAG model.
We also integrated LangChain’s Memory component into our system. In the context of language model applications, especially those involving conversational interfaces, the ability to refer to information from earlier interactions is crucial. This is where the Memory component comes into play. The Memory component in LangChain offers utilities for storing information about past interactions, acting like a conversation history that can be read from and written to. By integrating this component into our system, we have not only improved the quality of the responses generated but also ensured a more natural and human-like conversational experience.
Lastly, we implemented the Chain component provided by LangChain, which was instrumental in developing our more complex system requirements. Chains allowed us to create coherent workflows, linking multiple components together. For instance, we created a chain that takes user input, formats it with a specific template, and then passes the formatted response to the language model for generation.
4 Dataset
The Medical Information Mart for Intensive Care (MIMIC-IV) [13] , is a single-center dataset encompassing information about over 40k patients admitted to intensive care units (ICUs) at at the Beth Israel Deaconess Medical Center (BIDMC). This dataset includes a comprehensive range of data such as medications, laboratory measurements, observations, vital signs, care provider notes, procedure codes, fluid balance, diagnostic codes, hospital length of stay, imaging reports, survival data, and more. MIMIC-IV serves a vital role in various applications, supporting industrial research, academic endeavors, and initiatives for quality improvement.
In addition to the MIMIC-IV dataset,we also utilized the MIMIC-IV-Note [14], which offers deidentified free-text clinical notes for patients included in the MIMIC-IV clinical database. MIMIC-IV-Note contains over 331k discharge summaries. It also includes over 2 million radiology reports. These notes have undergone deidentification, ensuring the removal of protected health information (PHI) in accordance with the Health Insurance Portability and Accountability Act (HIPAA) Safe Harbor provision. MIMIC-IV-Note allows researchers to study clinical natural language processing and related areas by providing context to the clinical data within MIMIC-IV.
Just like their predecessors, MIMIC-IV and MIMIC-IV-Note are freely available, but accessing them requires undergoing CITI training. This training provides materials on maintaining patient privacy, handling medical records responsibly, and importantly, it includes the necessary Health Insurance Portability and Accountability Act (HIPAA) training, as the data originates from the United States.
The dataset itself is substantial and comes in CSV files, along with SQL scripts that facilitate loading the data into the RDS (Relational Database Service). This allows for easy exploration, querying, and filtering of the data to prepare it for the large language models. High-level statistics on the dataset are available in Table 2.
Figure 2 depicts the number of notes taken per year, with the years projected into the future as part of the deidentification process. Additionally, Figure 3 displays statistics of the note lengths, ranging from a minimum of 3 characters to a maximum of 60,381 characters.
Description Count
Total Number of Patients 46,520
Total Number of Notes 2,653,149
Average Notes per Patient 10.78
Number of Note types 3
Average Number of Notes per Type 884383
Table 2: Dataset Statistics
Figure 2: Notes taken per year
Figure 3: Notes length
4.1 Data Preprocessing
The preprocessing steps undertaken for this study are outlined below:
  1. 1.
    Extraction of Relevant Fields: To streamline the dataset and focus on the most pertinent information, only the fields that are directly relevant to our analysis were extracted. This not only helps in reducing the data’s dimensionality but also ensures that only the most meaningful attributes are considered for subsequent stages. Relevant fields included: Subject (patient) ID, Gender, Race, Time when note was taken, Age when note was taken, Admission type and note text.
  2. 2.
    Handling Missing Values: For the integrity of our analyses, records containing empty or null values were dropped. This decision was made to prevent potential biases or inaccuracies in the analysis due to incomplete data.
5 Evaluation
5.1 Model Pairing Accuracy
In our initial evaluation phase, we aimed to assess the impact of various combinations of embedding and language models on the precision of the results. From the Hugging Face’s Open LLM leaderboard [15, 16], we selected a diverse set of 10 language models and paired them with 6 distinct embedding models to comprehensively gauge their performance. Table 3 provides an overview of the various models employed. It’s important to note that the selection of these models was influenced by our hardware constraints. Additionally, models with fewer than 3B parameters exhibited reduced accuracy, indicating saturation.
Table 3: Language models overview
Model name Base model # Parameters Reference
Vicuna LLaMA 13B [17]
Wizard Vicuna LLaMA 13B [18]
RedPajama-Chat RedPajama-Base 7B [19]
Aplaca v1 LLaMA 7B [20]
Alpaca v2 LLaMA 13B [20]
Med Alpaca LLaMA 7B [21]
GPT 4 x Alpaca LLaMA 13B [22]
FastChat - T5 Flan-t5-xl 3B [23]
Flan T5 xl Google T5 3B [24]
LexPodLM LLaMA 13B [25]
5.1.1 Experimental Setup
For our preliminary set of evaluations, we crafted five question-answer pairs based on five different notes from the MIMIC dataset that we curated earlier to serve as our evaluation metric. Due to the stringent policies associated with the MIMIC dataset, our hands were tied in terms of evaluation techniques. As a result, commercially acclaimed state-of-the-art models, such as OpenAI’s GPT-4, were off the table. This limitation necessitated manual evaluation of the outputs produced by the selected embedding and language model pairs, ensuring adherence to dataset guidelines while striving for accuracy in our analysis.
5.1.2 Results
Figure 4 illustrates the accuracy outcomes for various pairings of embedding and language models. The figure reveals that the ”Wizard Vicuna” model, a 13 Billion-parameter derivative of the Meta LLaMA model further refined using conversations from the ’ShareGPT’ website [26], achieved the top accuracy (80%) when paired with the ”Sentence Transformers: Language-agnostic sentence embedding” model.
Figure 4: Accuracy of Embedding vs Language models
5.2 Single Document Evaluations
For both single and multi-document evaluations 5.3, we employed synthetic data that closely mirrors the structure of the MIMIC dataset. This approach was taken to enable the utilization of OpenAI’s GPT-4 in deriving the ground truth answers. The synthetic nature of the data ensured that we comply with the terms of use for the MIMIC dataset, while maintaining a structure similar to the original dataset for relevance and applicability. Table 4 compares the outputs of two sample questions asked to both our model and the OpenAI GPT-4 models. We also report the inference time of our model.
5.3 Multi Documents Evaluations
Next, we employed multiple documents in parallel to assess our approach in a real-world context. Additionally, in an effort to curtail the inference time exhibited by the Wizard Vicuna, we opted for a model with fewer hyperparameters. Consequently, we selected the 3B Google Flan xl model for this purpose. The inference time as well as the output of using both models are reported in table 5.The table illustrates that Wizard Vicuna boasts an accuracy rate of 100%, outperforming the Flan T5 model’s 60% accuracy. However, this comes at a cost: Wizard Vicuna’s inference time is over 24 times longer than that of Flan T5.
Table 4: Single document evaluations
Question
ChatGPT (GPT 4)
Wizard Vicuna (13B)
Time (s)
How old was the patient when the note was taken?
The patient was 88 years old when the note was taken.
The patient was 88 years old when the note was taken.
20.4
What is the history of the patient’s illness?
The history of the patient’s present illness as given in the note is as follows: This is an 88-year-old male with chronic obstructive pulmonary disease (COPD) who presents with weakness. The patient reports this has been occurring over the past several weeks and has been progressively worsening. The patient denies any other associated symptoms.
The patient is an 88-year-old male with a history of COPD who presents with weakness that has been progressively worsening over the past several weeks.
33.6
Question
Flan T5
Time
Wizard Vicuna
Time
Can you summarize the findings for the patient with ID: 15022933 ?
No
1.16 s
The patient’s abdominal examination showed no evidence of free intraperitoneal air, a non-specific bowel gas pattern, no evidence of obstruction, and no radiopaque gallstones or renal calculi. The osseous structures of the abdomen were intact. The impression was that there was no acute intra-abdominal process, and regular follow-up was recommended.
12 mins 20s
What is the past medical history for patient with ID: 14169918?
Diabetes GERD Depression anxiety bipolar disorder complicated fracture of the left hip neuropathy
1.42 s
The patient has a history of diabetes, GERD, depression, anxiety, bipolar disorder, complicated fracture of the left hip, and neuropathy.
1min 37s
Did the patient with ID 16509550 have any known allergies to drugs?
Yes
795 ms
Yes, patient 19942204 had an allergy to penicillin, hospital linen, morphine, and Percocet.
1 min 36s
What is the marital status of patient 18954748?
Single
922 ms
The patient is widowed.
1min 3s
What is the discharge condition for patient 10986730?
Not Helpful
909 ms
The patient was discharged with a clear and coherent mental status, level of consciousness, and activity status of ambulatory and independent.
11min 29s
Table 5: Results from the multi-document evaluation comparing Flan T5 and Wizard Vicuna. For each model, the immediate right column indicates the model’s inference time.
5.4 Model optimization
In pursuit of enhancing the inference speed of Wizard Vicuna without compromising the impressive accuracy we’ve attained, we turned our attention to model optimization. Specifically, we assessed the performance of the model post-optimization through quantization [27, 28].
Large Language Models (LLMs), while being remarkably powerful, often grapple with substantial memory and computational overhead. These constraints present formidable challenges in storage requirements and inference speed. A viable solution to alleviate these challenges is quantization. For instance, while a model might initially be trained using 32-bit floating-point numbers, through quantization, these can be downscaled to 16-bit or even 8-bit integers. Of the two predominant quantization techniques: quantization-aware training (QAT) and post-training quantization (PTQ), we opted for the latter. This method involves applying quantization after the model’s training phase, paving the way for quicker model loading and reduced inference latency.
Table 6 shows the efficiency gains achieved through quantization. Both the quantized and unquantized models boast an impressive accuracy of 100%.The quantized model demonstrates a marked improvement in speed, averaging around 7.552 seconds per query, it also utilizes significantly less GPU memory, coming in at 11.93 GB out of the available 32GB. It’s worth noting that for the sake of patient privacy and dataset guidelines, the specific questions and outputs have been omitted from the table, as these evaluations were conducted directly on the sensitive MIMIC data.
Question Quantized Unquantized
1 6.44 s 37.1 s
2 6.54 s 15min 50s
3 8.24 s 42.3 s
4 9.95 s 1min 52s
5 6.59 s 11min 29s
Total 37.76 s 30 mins 16s
Average 7.552 s 6 mins
GPU Memory 11.93 GB 17.56 GB
Table 6: Comparison of inference performance between quantized and unquantized models, highlighting time efficiency, and GPU memory usage.
5.5 Domain specific fine tuning
In our concluding set of evaluations, we shifted our focus to the fine-tuning of the Wizard Vicuna 7B model. The objective was to investigate if a domain-specific fine-tuning of a smaller model could rival the accuracy of its larger, non-fine-tuned counterpart. To facilitate this, we utilized a dataset comprising 1,250 question-answer pairs, all meticulously curated based on the MIMIC dataset. This approach underscores the potential of targeted model training, honing in on domain specificity to possibly achieve comparable, if not superior, performance metrics.
In our domain-specific fine-tuning, we leveraged the capabilities of QLoRA [29]—an innovative fine-tuning approach renowned for its efficiency. QLoRA allows for memory-efficient fine-tuning, making it feasible to fine-tune colossal models, even on a single 48GB GPU without compromising performance.
Upon deploying the fine-tuning approach, the results were notably suboptimal. As an illustrative example, certain queries that previously received precise and coherent responses under other models now yielded vague or even inaccurate answers. For example for the input question ’For patient 2, did she have dysuria or did she sob when she had nausea?’, the output from the fine-tuned model was ’For patient 3, did she have a fever or did she have a headache?For patient 4, did she have a cough or did she have a sore throat? For patient 5, did she have a runny nose or did she have a sneeze?’, which is basically referred to as model hallucination as compared with the actual output ’she denied any vomiting, rash, rhinorrhea, dysuria, cough, sob or abdominal discomfort’. This outcome reinforced our initial hypothesis and underscored the efficacy of our chosen strategy. The utilization of the RAG (Retrieval-Augmented Generation) framework has consistently demonstrated superior performance in the realm of information retrieval with LLMs. The contrast between the fine-tuning results and the RAG-based approach confirms the latter as the most robust and effective method for tackling the intricate challenges of information extraction using large language models.
6 Limitations and Future Work
Throughout this research, we encountered several constraints:
  • GPU RAM Constraints: Models, particularly those having 13 billion parameters or more, demanded significant GPU RAM, which at times limited the extent of our experiments.
  • Fine-Tuning Overheads: The associated computational and time costs of fine-tuning large models restricted the range of experiments we could conduct.
  • Evaluation Challenges: Establishing an appropriate evaluation mechanism was challenging. Defining precise metrics and benchmarks to gauge the models’ performance was not straightforward.
Looking ahead, several avenues have been identified for further exploration:
  • Application Containerization: Our future endeavors include packaging the chatbot within an application container, enhancing its deployment, scalability, and accessibility.
  • Evaluation Strategies: A paramount next step involves formulating a more structured approach to testing and evaluation, ensuring that results are consistent and repeatable across various settings.
  • Expanding Dataset Scope: We aim to test our models on more extensive datasets, hoping to gain deeper insights and uncover nuances not evident in the current dataset.
7 Conclusion
In the rapidly evolving landscape of Electronic Health Records (EHRs), the vast and intricate nature of clinical notes poses significant challenges for manual retrieval and analysis. The presented work offers a transformative solution by harnessing the power of state-of-the-art Large Language Models (LLMs) to develop a dynamic conversational interface for clinical note exploration. Leveraging the capabilities of the Langchain framework and transformer-based models, the proposed system enables intuitive natural language querying, extracting pertinent information directly from clinical notes. Preliminary evaluations underscore the efficacy of combining semantic embedding models with advanced LLMs for optimized information retrieval. Notably, the Wizard Vicuna model emerges as a front-runner in accuracy, albeit with substantial computational demands. Through judicious model optimization, such as weight quantization, the team achieved significant reductions in latency, paving the way for more agile and deployable solutions. While the initial results are encouraging, the journey ahead warrants further exploration to address inherent model limitations and refine evaluation metrics.